OpenStudy (anonymous):

Diane has quarters pennies and dimes, her total is 1.55, there are 4 more dimes than quarters and the number of pennies equals the sums of the number of quarters and dimes, how many pennies does she have? I got one of the equations of 35q+1p=115 but dont know where to go from there

OpenStudy (larseighner):

It is always easier to work the problems with whole cents instead of dollars and fractions, so 1p + 10d + 25q = 155 q+4 = d p = d + q If you find p in terms of q using the latter 2 equations, you can get from where you are to the solution.

OpenStudy (anonymous):

Got it thanks, It is 10 pennies!